# Self-Adaptive Physics-Informed Quantum Machine Learning for Solving Differential Equations

### v2

@@ Abstract
Chebyshev polynomials have shown significant promise as an efficient tool for both classical and quantum neural networks to solve linear and nonlinear differential equations. In this work, we adapt and generalize this framework in a quantum machine learning setting for a variety of problems, including the 2D Poisson’s equation, second-order differential equation, system of differential equations, and nonlinear Riccati equation. In particular, we propose in the quantum setting a modified Self-Adaptive Physics-Informed Neural Network (SAPINN) approach, where self-adaptive weights are applied to problems with multi-objective loss functions. We further explore capturing correlations in our loss function using a quantum-correlated measurement, resulting in improved accuracy for initial value problems. We analyse also the use of entangling layers and their impact on the solution accuracy for second-order differential equations. The results indicate a promising approach to the near-term evaluation of differential equations on quantum devices.
@@ 1 Introduction
In our work, we combine the approaches of self-adaptive PINNs with quantum feature maps. We extend the usage of the quantum Chebyshev feature map while solving a variety of differential equations. Our results show that combining these approaches leads to a measurable improvement in the convergence and accuracy for our test cases, namely the Riccati equation, system of differential equations, second-order differential equation, and the 2D Poisson’s equation. Particularly, we incorporate the approaches of self-adaptive PINNs via variational quantum circuits for balancing multi-objective loss functions. Moreover, we study the use of correlated measurement observables and more entangling layers for better convergence. Our results show that these approaches lead to a measurable improvement in the convergence and accuracy for solving the given test cases, namely the Riccati equation, system of differential equations, second-order differential equation, and the 2D Poisson’s equation.
@@ 2 Methods
The variational quantum circuit to solve differential equations first proposed by Kyriienko et al. [21] is presented in Fig. 1. In this circuit, there are three main parts, namely, a quantum feature map circuit for encoding Chebyshev polynomials, a variational circuit optimizing the cost function, and measurement (where we test correlated observables in the quantum loss functions).
@@ 3 Results
For this work, we investigate and benchmark our proposed methodology on selected problems, based on earlier literature [21, 24]. We start with a highly nontrivial dynamics problem such as the Riccati equation which is quadratic in unknown function. Then we consider a coupled first-order differential equation, formulated as a system of DEs. In addition, a second-order differential equation is investigated. Finally, a problem on the two-dimensional Poisson’s equation presented by Kharazmi et al. [24] is analyzed. All simulations are performed using Pennylane software [25].
@@ 3.1 Complex Dynamics of Riccati Equation
This DE is first solved by two classical methods: first, numerically by the Runge-Kutta method, which is considered the true solution for further comparisons, and second, using PINNs, which is considered as the classical ML solution. In the case of PINNs, a two-layer feed-forward neural network with 32 hidden neurons was considered, which has a $\tanh$ activation function in each layer. To solve this problem using PINNs, the loss from initial conditions and differential equations is simply added as a multi-objective loss function A. The results are visualized in Fig. 3(a). We first carried out the optimization with the Adam optimizer and found that it was unable to converge to the desired solution. However, when we carried out the optimization with the L-BFGS algorithm, the desired solution was reached.
For the quantum solution, a quantum circuit with 8 qubits and a depth of 8 blocks in the variational circuit with a circular entanglement layout (refer to B.2) has been considered. The differential equation has terms proportional to $u$ and $u^{2}$, which enables incorporating the initial condition into floating boundary handling (refer to B.5). This enables faster convergence since a single combined loss function is used instead of a slower multi-objective optimization [21]. A set of $N_{f}=100$ collocation points with $x_{f}^{i}\in\{1,\cdots,N_{f}\}$ are generated by equidistant grid points. The initial condition is given at $x_{f}^{1}$. Consider the circuit output as $u_{\text{pred}}$, and we define the modified output $u_{\text{m}}$ such that
@@
Figure 4: Solving of a system of coupled differential equations (refer Eq. 19, Eq. 20, and Eq. 21). Visualization of (4(a)) comparison between true solution obtained classically via the Runge-Kutta method as well as the quantum solution for $u_{1}(x)$ and $u_{2}(x)$ as well as (4(b)) the loss for the summation and tensor product of Pauli-Z operators cases over the number of optimization iterations.
@@
Thus, the Riccati equation is a promising first example, where quantum PINNs can effectively encode the solutions, especially in systems where fast multi-partite encoding can be obtained before the measurement.
Figure 5: Solution of second-order DE (refer Eq. 22) using constant weights for balancing multi-objective loss function. Visualization of (5(a)) reference variational quantum circuit (refer Fig. 1 and Ref. [21]) (referred as RC). It is arranged as RZ, RX, RZ rotational layers followed by CNOT gates in cyclical order. (5(b)) Modified variational circuit with added entangling layers in between rotational layers (referred to as AEC). (5(c)) Solutions $u(x)$ obtained from three cases and their comparison with the true solution. Case 1 corresponds to employing of RC variational circuit, and balancing weights for multi-objective loss function $\alpha_{f}=\alpha_{b}=1$. Case 2 corresponds to the RC variational circuit, and weights $\alpha_{f}=10^{-1}$ and $\alpha_{b}=10^{3}$. Case 3 corresponds to the AEC variational circuit, and weights $\alpha_{f}=10^{-1}$ and $\alpha_{b}=10^{3}$. (5(d)) Comparison of accuracy in terms of $1-R^{2}$ measurement over iterations among three cases, and checking of (5(e)) first derivative $\frac{du}{dx}$ and (5(f)) second derivative $\frac{d^{2}u}{dx^{2}}$ of solution after training.
@@ 3.2 System of Differential Equations
We now turn our methodology to another problem, first proposed in Kyriienko et al. [21], where a first-order system of differential equations is investigated. Once again, we use the same quantum PINN method as before. In this case, a vector $\boldsymbol{u}(x)=\begin{bmatrix}u_{1}(x),u_{2}(x)\end{bmatrix}$ describes two modes, whose progression in time is $x$ given by
$F_{1}\left(\mathop{}\!\mathrm{d}_{x}\boldsymbol{u},\boldsymbol{u},x\right)=$ $\frac{\mathop{}\!\mathrm{d}u_{1}}{\mathop{}\!\mathrm{d}x}-\lambda_{1}u_{2}-\lambda_{2}u_{1}=0\,,$ (19)
$F_{2}[\mathop{}\!\mathrm{d}_{x}\boldsymbol{u},\boldsymbol{u},x]=$ $\frac{\mathop{}\!\mathrm{d}u_{2}}{\mathop{}\!\mathrm{d}x}+\lambda_{2}u_{2}+\lambda_{1}u_{1}=0\,,$ (20)
@@
where $\lambda_{1}$ and $\lambda_{2}$ are coupling parameters and $u_{1,0}$ as well as $u_{2,0}$ are initial conditions. To simulate this problem, a quantum circuit with three qubits and a depth of seven blocks in the variational circuit is considered (see B.2). An Adam optimizer is once again chosen, with a learning rate of 0.1. The observables are compared for the cases of addition and product as explained in the previous example. The domain is considered to be $x\in[0,0.9]$, discretized into 100 equidistant grid points. The parameters for this problem are $\lambda_{1}=5,\,\lambda_{2}=3$ and initial conditions are $u_{1,0}=0.5$ and $u_{2,0}=0$.
The results obtained from global observable are shown in Fig. 4(a). It is clear that quantum PINN circuits accurately capture the known solution. In Fig. 4(b), we again compare the modified observables such as the tensor product with the summation of Pauli-Z operators. Once again we see a large impact in the accuracy and convergence towards the known solution.
3.3 Second-Order Differential Equation
@@
Figure 7: Solution of 2D Poisson’s equation (refer Eq. 25). Visualization of (7(a)) True and (7(b)) predicted solution $u(x,y)$ of Poisson’s equation with color bar mapping the function values $u(x,y)$ to contour plot. Interpolation test for 20 and 100 equidistant points at (7(c)) $x=0.25$ and (7(d)) $y=0.75$, where the problem is originally solved at 30 equidistant points. 7(e)) The accuracy comparison in terms of $1-R^{2}$ measurement over iterations for reference (RC) and added entanglement circuit (AEC).
@@
The differential equation in this problem does not contain the terms $u(x)$ in it. Therefore, the Dirichlet boundary conditions cannot be applied using floating boundary handling (refer B.5). Hence, the loss function is formulated as a multi-objective loss function (A). This loss function requires balancing the loss terms $\alpha_{f},\alpha_{b}$. Therefore, two cases are compared: in the first case, we show the effect of unbalanced loss terms, where the weights of the total loss function are given with $\alpha_{f}=\alpha_{b}=1$ (Case 1), and in the second case, the weights are constant factors and they are heuristically chosen as $\alpha_{f}=10^{-1}$ and $\alpha_{b}=10^{3}$ (Case 2). For these cases, a circuit with three qubits and a depth of seven blocks in the variational circuit B.2 has been considered where the observables are chosen with the standard $\hat{\mathcal{C}}=\sum_{j}Z_{j}$. The obtained solution for all three cases together with the first and second derivative of the solution is visualized in Fig. 5(c), Fig. 5(e), and Fig. 5(f). The accuracy of the solutions is compared through the $R^{2}$ score using the term $1-R^{2}$ (see Fig. 5(d)). The results comparison between cases 1 and 2 suggest that balancing weights are important while handling multi-objective loss function. However, the accuracy of the quantum ML circuit has trouble in stabilizing below 1%.
@@ 3.4 Second-Order Partial Differential Equation
The quantum formulation of this problem is described in detail in Section 2.4. The cost function for this problem consists of residual loss and boundary conditions. For calculating the residual loss, the domain for this problem is discretized into $30\times 30$ equidistant points totaling 900 collocation points $N_{f}(x_{f},y_{f})$. The solution on the boundary is represented by $N_{b}$ labeled data points $\{x_{b}^{i},y_{b}^{i},u_{b}^{i}\}$ and the Dirichlet boundary conditions are given by $u_{b}(x,0),u_{b}(x,0.9),u_{b}(0,y),$ and $u_{b}(0.9,y)$. Note that, we are examining a slightly altered problem domain than the usual boundary points $x=1,y=1$ due to computational costs and numerical instabilities in the differential operator leading to a modified boundary (see Section 2.1). Finally, the cost function is formulated as $C=\alpha_{f}\text{MSE}_{f}+\alpha_{b}\text{MSE}_{b}$ (see Appendix A), where the weights $\alpha_{f}=10^{-1}$ and $\alpha_{b}=10^{3}$. Adam optimizer with a learning rate of $0.1$ is employed with 300 iterations.

### v3

@@ Abstract
Chebyshev polynomials have shown significant promise as an efficient tool for both classical and quantum neural networks to solve linear and nonlinear differential equations. In this work, we adapt and generalize this framework in a quantum machine learning setting for a variety of problems, including the 2D Poisson’s equation, second-order linear differential equation, system of differential equations, nonlinear Duffing and Riccati equation. In particular, we propose in the quantum setting a modified Self-Adaptive Physics-Informed Neural Network (SAPINN) approach, where self-adaptive weights are applied to problems with multi-objective loss functions. We further explore capturing correlations in our loss function using a quantum-correlated measurement, resulting in improved accuracy for initial value problems. We analyse also the use of entangling layers and their impact on the solution accuracy for second-order differential equations. The results indicate a promising approach to the near-term evaluation of differential equations on quantum devices.
@@ 1 Introduction
In our work, we extend the usage of the quantum Chebyshev feature map while solving a variety of differential equations. In addition, we incorporate the approach of self-adaptive weights for balancing multi-objective loss functions while solving differential equations. Particularly, we modify the method of self-adaptive weights introduced for classical PINNs [23] into the setting of the proposed variational quantum algorithm. Moreover, we study the use of correlated measurement observables and more entangling layers for better convergence. Our results show that these approaches lead to a measurable improvement in the convergence and accuracy for solving the given test cases, namely the Duffing equation, Riccati equation, system of differential equations, second-order linear differential equation, and the 2D Poisson’s equation.
@@ 2 Methods
The variational quantum circuit to solve differential equations proposed by Kyriienko et al. [21] is presented in Fig. 1. In this circuit, there are three main parts, namely, a quantum feature map circuit for encoding Chebyshev polynomials, a variational circuit optimizing the cost function, and measurement (where we test correlated observables in the quantum loss functions). An overview of the variables used in this work is mentioned in Table 1 for the sake of clarity.
Table 1: Overview of used variables.
Variables
$N_{f}$ Set of collocation points
$N_{b}$ Set of boundary points
$x_{f}^{i}$ Collocation grid points
$x_{b}^{i}$ Boundary points
$u$/$u_{\text{true}}$ True solution
$u_{\text{pred}}$ Predicted solution
$\alpha_{f}$ Residual loss weight
$\alpha_{b}$ Boundary loss weight
$\hat{\mathcal{U}}_{\phi}$ Quantum feature map
$\hat{\mathcal{U}}_{\theta}$ Variational quantum circuit
@@ 3 Results
For this work, we investigate and benchmark our proposed methodology on selected problems, based on earlier literature [21, 24]. We start with a highly nontrivial dynamics problem such as the Riccati equation which is quadratic in an unknown function. Then we consider a coupled first-order differential equation, formulated as a system of DEs. In addition, a second-order linear differential equation and the nonlinear Duffing equation are investigated. Finally, a problem on the two-dimensional Poisson’s equation presented by Kharazmi et al. [24] is analyzed. All simulations are performed using Pennylane software [25].
@@ 3.1 Complex Dynamics of Riccati Equation
This DE is first solved by two classical methods: first, numerically by the explicit Runge-Kutta method of third order [26], which is considered the true solution for further comparisons, and second, using PINNs, which is considered as the classical ML solution. In the case of PINNs, a two-layer feed-forward neural network with 32 hidden neurons was considered, which has a $\tanh$ activation function in each layer. To solve this problem using PINNs, the loss from initial conditions and differential equations is simply added as a multi-objective loss function A. The results are visualized in Fig. 3(a). We first carried out the optimization with the Adam optimizer and found that it was unable to converge to the desired solution. However, when we carried out the optimization with the L-BFGS algorithm, the desired solution was reached.
For the quantum solution, a quantum circuit with 8 qubits and a depth of 8 blocks in the variational circuit with a circular entanglement layout (refer to B.2) has been considered. The differential equation has terms proportional to $u$ and $u^{2}$, which enables incorporating the initial condition into floating boundary handling (refer to B.5). This enables faster convergence since a single combined loss function is used instead of a slower multi-objective optimization [21].
A set of $N_{f}=100$ collocation points with $x_{f}^{i}\in\{1,\cdots,N_{f}\}$ are generated by equidistant grid points. The initial condition is given at $x_{f}^{1}$. Consider the circuit output as $u_{\text{pred}}$, and we define the modified output $u_{\text{m}}$ such that
@@
Figure 4: Solving of a system of coupled differential equations (refer Eq. 19, Eq. 20, and Eq. 21). Visualization of (4(a)) comparison between classical and quantum solution for $u_{1}(x)$ and $u_{2}(x)$ as well as (4(b)) the loss for the summation and tensor product of Pauli-Z operators cases over the number of optimization iterations.
@@
Thus, the Riccati equation is a promising first example, where the quantum approach can effectively encode the solutions, especially in systems where fast multi-partite encoding can be obtained before the measurement.
Figure 5: Solution of second-order DE (refer Eq. 22) using constant weights for balancing multi-objective loss function. Visualization of (5(a)) reference variational quantum circuit (refer Fig. 1) (referred to as RC). It is arranged as RZ, RX, RZ rotational layers followed by CNOT gates in cyclical order. (5(b)) Modified variational circuit with added entangling layers in between rotational layers (referred to as AEC). (5(c)) Solutions $u(x)$ obtained from three cases and their comparison with the true solution. Case 1 corresponds to employing of RC variational circuit, and balancing weights for multi-objective loss function $\alpha_{f}=\alpha_{b}=1$. Case 2 corresponds to the RC variational circuit, and weights $\alpha_{f}=10^{-1}$ and $\alpha_{b}=10^{3}$. Case 3 corresponds to the AEC variational circuit, and weights $\alpha_{f}=10^{-1}$ and $\alpha_{b}=10^{3}$. (5(d)) Comparison of accuracy in terms of $1-R^{2}$ measurement over iterations among three cases, and checking of (5(e)) first derivative $\frac{du}{dx}$ and (5(f)) second derivative $\frac{d^{2}u}{dx^{2}}$ of solution after training.
@@ 3.2 System of Differential Equations
We now turn our methodology to another problem, first proposed in Kyriienko et al. [21], where a first-order system of differential equations is investigated. In this case, a vector $\boldsymbol{u}(x)=\begin{bmatrix}u_{1}(x),u_{2}(x)\end{bmatrix}$ describes two modes, whose progression in time is $x$ given by
$F_{1}\left(\mathop{}\!\mathrm{d}_{x}\boldsymbol{u},\boldsymbol{u},x\right)=$ $\frac{\mathop{}\!\mathrm{d}u_{1}}{\mathop{}\!\mathrm{d}x}-a_{1}u_{2}-a_{2}u_{1}=0\,,$ (19)
$F_{2}[\mathop{}\!\mathrm{d}_{x}\boldsymbol{u},\boldsymbol{u},x]=$ $\frac{\mathop{}\!\mathrm{d}u_{2}}{\mathop{}\!\mathrm{d}x}+a_{2}u_{2}+a_{1}u_{1}=0\,,$ (20)
@@
where $a_{1}$ and $a_{2}$ are coupling parameters and $u_{1,0}$ as well as $u_{2,0}$ are initial conditions. The classical solution for this problem is obtained through ODE solvers by [27, 28]. To achieve a quantum solution, a quantum circuit comprising three qubits and a depth of seven blocks in the variational circuit is considered (see B.2). An Adam optimizer is once again chosen, with a learning rate of 0.1. The observables are compared for the cases of addition and product as explained in the previous example. The domain is considered to be $x\in[0,0.9]$, discretized into 100 equidistant grid points. The parameters for this problem are $a_{1}=5,\,a_{2}=3$ and initial conditions are $u_{1,0}=0.5$ and $u_{2,0}=0$.
Figure 7: Solution of Duffing equation (refer Eq. 25). Visualization of (7(a)) displacement $(u)$ vs. time $(t)$, (7(b)) velocity $(\frac{\mathop{}\!\mathrm{d}u}{\mathop{}\!\mathrm{d}t})$ vs. time $(t)$, (7(c)) velocity vs. displacement over the time evolution, also known as the phase-space plot of Duffing oscillation. The color bar on the right depicts the time $(t)$. The accuracy comparison (7(d)) in terms of $1-R^{2}$ measurement over iterations for reference (RC) and added entanglement circuit (AEC).
The results obtained from global observable are shown in Fig. 4(a). It is clear that the proposed quantum algorithm accurately captures the known solution. In Fig. 4(b), we again compare the modified observables such as the tensor product with the summation of Pauli-Z operators. Once again we see a large impact in the accuracy and convergence towards the known solution.
3.3 Second-Order Linear Differential Equation
@@
The differential equation in this problem does not contain the terms $u(x)$ in it. Therefore, the Dirichlet boundary conditions cannot be applied using floating boundary handling (refer B.5). Hence, the loss function is formulated as a multi-objective loss function (A). This loss function requires balancing the loss terms $\alpha_{f},\alpha_{b}$. Therefore, two cases are compared: in the first case, we show the effect of unbalanced loss terms, where the weights of the total loss function are given with $\alpha_{f}=\alpha_{b}=1$ (Case 1), and in the second case, the weights are constant factors and they are heuristically chosen as $\alpha_{f}=10^{-1}$ and $\alpha_{b}=10^{3}$ (Case 2). For these cases, a circuit with three qubits and a depth of seven blocks in the variational circuit B.2 has been considered where the observables are chosen with the standard $\hat{\mathcal{C}}=\sum_{j}Z_{j}$.
The obtained solution for cases 1 and 2 together with the first and second derivative of the solution is visualized in Fig. 5(c), Fig. 5(e), and Fig. 5(f). The accuracy of the solutions is compared through the $R^{2}$ score using the term $1-R^{2}$ (see Fig. 5(d)). The results comparison between cases 1 and 2 suggest that balancing weights are important while handling multi-objective loss function. However, the accuracy of the quantum ML circuit has trouble in stabilizing below 1%.
@@
3.4 Second-Order Nonlinear Differential Equation
For the next problem, we consider a Duffing equation which is a nonlinear second-order differential equation given by,
$\frac{\mathop{}\!\mathrm{d}^{2}u}{\mathop{}\!\mathrm{d}t^{2}}+\delta\frac{\mathop{}\!\mathrm{d}u}{\mathop{}\!\mathrm{d}t}+\alpha u+\beta u^{3}=\gamma\cos{(\omega t)}.$ (25)
In this work, we have considered an example of this problem from [33], where the parameters are given by, $\alpha=1,\beta=1,\delta=0.1,\gamma=0.1,\omega=0.4$. The initial conditions are given by $u(0)=0$ and $\frac{\mathop{}\!\mathrm{d}u(0)}{\mathop{}\!\mathrm{d}t}=0$. The time evolution to solve the problem is chosen from $t=0$ to $t=20$. As a reference solution (also referred to as true solution), we employed an explicit Runge-Kutta method of fifth order [34]. For the quantum solution, the dependent variable is suggested to be not close to 1.0 due to singularity (refer Section 2.1). Therefore, the dependent variable in the Duffing equation is scaled and solved within the interval $t=[0,l]$, where we choose the limit, $l=0.9$. Consider the total time of the evolution as $T=20$ and the scaled time variable as $\tau=\frac{t}{T}l$. Then the modified differential equation is given by,
$\left(\frac{l}{T}\right)^{2}\frac{\mathop{}\!\mathrm{d}^{2}u}{\mathop{}\!\mathrm{d}\tau^{2}}+\frac{l}{T}\delta\frac{\mathop{}\!\mathrm{d}u}{\mathop{}\!\mathrm{d}\tau}+\alpha u+\beta u^{3}=\gamma\cos{\left(\frac{\omega T\tau}{l}\right)}.$ (26)
The solution grid is chosen to be the discretization of 50 equidistant time intervals. Similar to the previous problem, the cost function is formulated as $C=\alpha_{f}\text{MSE}_{f}+\alpha_{b}\text{MSE}_{b}$ (see A), where the weights $\alpha_{f}=1$ and $\alpha_{b}=1$. In this case, there was no need for balancing these two terms while their magnitude was found to be of similar order. The hyperparameters chosen for this problem are five qubits and the depth of eight blocks in the variational part of the quantum circuit (refer to B.2). The observables are chosen to be $\hat{\mathcal{C}}=\sum_{j}Z_{j}$. To test the applicability of the proposed added entangling layers, we have considered two cases similar to the previous problem. The first case is where we chose the reference circuit Fig. 5(a) and the second one is the added entangling circuit Fig. 5(b). For both cases, the initial variational parameters are the same for fair comparison. An Adam optimizer with a learning rate of 0.01 is employed with 300 iterations. After solving the problem at coarse 50 discretized points, the trained circuit is tested to predict the solution at 500 equidistant time steps for visualization of a smooth solution. This is one of the advantages of the spectral methods over finite difference methods, namely of having smooth interpolation. The results presented in the Fig. 7(a) and Fig. 7(b) show that the displacement and velocity are well-matched with the Runge-Kutta solution. We have also depicted the phase space plot in Fig. 7(c), which is velocity vs. displacement, to visualize the resulting Duffing oscillation. Furthermore, the accuracy plot using the measurement $1-R^{2}$ (see Fig. 7(d)) clearly shows that the added entangling circuit has better performance as compared to the reference circuit.
Figure 8: Solution of 2D Poisson’s equation (refer Eq. 27). Visualization of (8(a)) true and (8(b)) predicted solution $u(x,y)$ of Poisson’s equation with color bar mapping the function values $u(x,y)$ to contour plot. Interpolation test for 20 and 100 equidistant points at (8(c)) $x=0.25$ and (8(d)) $y=0.75$, where the problem is originally solved at 30 equidistant points. 8(e)) The accuracy comparison in terms of $1-R^{2}$ measurement over iterations for reference (RC) and added entanglement circuit (AEC).
@@ 3.5 Second-Order Partial Differential Equation
The quantum formulation of this problem is described in detail in Section 2.4. The cost function for this problem consists of residual loss and boundary conditions. For calculating the residual loss, the domain for this problem is discretized into $30\times 30$ equidistant points totaling 900 collocation points $N_{f}(x_{f},y_{f})$. The solution on the boundary is represented by $N_{b}$ labeled data points $\{x_{b}^{i},y_{b}^{i},u_{b}^{i}\}$ and the Dirichlet boundary conditions are given by $u_{b}(x,0),u_{b}(x,0.9),u_{b}(0,y),$ and $u_{b}(0.9,y)$. Note that, we are examining a slightly altered problem domain than the usual boundary points $x=1,y=1$ due to computational costs and numerical instabilities in the differential operator leading to a modified boundary (see Section 2.1). Finally, the cost function is formulated as $C=\alpha_{f}\text{MSE}_{f}+\alpha_{b}\text{MSE}_{b}$ (see A), where the weights $\alpha_{f}=10^{-1}$ and $\alpha_{b}=10^{3}$. Adam optimizer with a learning rate of $0.1$ is employed with 300 iterations.